Find the equation of the line having given slope and containing the given point
m=7/8, (7,-9)

Answer by nerdybill(7384)

m=7/8, (7,-9)
plug into "point-slope" form:
y - y1 = m(x - x1)
y - (-9) = (7/8)(x - 7)
y + 9 = (7/8)x - 49/8
y = (7/8)x - 49/8 - 9
y = (7/8)x - 49/8 - 72/8
y = (7/8)x - 121/8 (this is what they're looking for)